# Both Text and Images Leaked! A Systematic Analysis of Data Contamination in Multimodal LLM

### v2

@@ 6.2 Tracing Origins: A Review of MLLM’s Visual Instruction Tuning Data
As Table 6 illustrates, MLLMs marked in red and yellow typically exhibit a significant contamination degree. Yet, even MLLMs labeled in green aren’t exempt from the risk of cross-modal contamination. This is because some models have been trained on large-scale interleaved image-text datasets (e.g., OBELICS (Laurençon et al., 2023)), datasets derived from online sources (e.g., Conceptual Caption (Sharma et al., 2018)), or in-house data. Furthermore, some models haven’t fully disclosed their training data, which may lead to overlooked potential leaks in benchmark datasets.
@@ 7 Conclusion and Future Work
In this study, we introduce and validate a multimodal data contamination detection framework, MM-Detect, providing new perspectives for evaluating contamination in MLLMs. We discovered that popular MLLMs exhibit varying degrees of data contamination, which directly impacts their performance and generalization ability. In addition, our experiment indicates that MM-Detect is sensitive to varying degrees of contamination and can highlight significant performance improvements due to the leakage in the multimodal benchmark training set. Furthermore, we found that the contamination in MLLMs may not solely originate from the cross-modal contamination but could also stem from the unimodal contamination.
@@
We acknowledge several limitations in our work. First, this work is limited to discussions around visual modalities, and does not yet cover other modalities such as audio or video. Second, we only selected widely used and representative multimodal datasets for detection, including multiple-choice datasets and caption datasets, without testing additional datasets, such as open-ended generation and cloze questions. However, we speculate that the method Slot Guessing for Perturbation Caption may also apply to other types of image-feature-analyzing benchmarks. Third, the effectiveness of Option Order Sensitivity Test can be undermined by option shuffling, which, while potentially improving model performance, is computationally expensive and may increase the training cost.

### v3

@@ 6.2 Analyzing Cross-modal Contamination in Multimodal Fine-tuning
As Table 6 illustrates, MLLMs marked in red and yellow typically exhibit a significant contamination degree. Yet, even MLLMs labeled in green aren’t exempt from the risk of cross-modal contamination. This is because some models have been trained on large-scale interleaved image-text datasets (e.g., OBELICS (Laurençon et al., 2023)), datasets derived from online sources (e.g., Conceptual Caption (Sharma et al., 2018)), or in-house data. Furthermore, some models haven’t fully disclosed their training data, which may lead to overlooked potential leaks in benchmark datasets. A detailed overlap analysis is in Appendix E.
@@ 7 Conclusion and Future Work
In this study, we systematically analyzed multimodal data contamination in MLLMs through our proposed detection framework, MM-Detect. We demonstrated that MM-Detect effectively quantifies and detects varying contamination degrees, revealing significant performance biases induced by benchmark leakage. Importantly, we identified that contamination originates notably from both unimodal pre-training and multimodal fine-tuning phases, impacting the reliability and fairness of multimodal evaluations.
@@
We acknowledge several limitations in our work. First, this work is limited to discussions around visual modalities, and does not yet cover other modalities such as audio or video. Second, we only selected widely used and representative multimodal datasets for detection, including multiple-choice datasets and caption datasets, without testing additional datasets, such as open-ended generation and cloze questions. However, we speculate that the method Slot Guessing for Perturbed Caption may also apply to other types of image-feature-analyzing benchmarks. Third, the effectiveness of Option Order Sensitivity Test can be undermined by option shuffling, which, while potentially improving model performance, is computationally expensive and may increase the training cost. Fourth, as a perturbation-based black-box detector, MM-Detect might underestimate contamination if a model generalizes sufficiently to answer perturbed questions correctly. Although dataset-level evaluations reduce this risk, completely eliminating such false-negative cases remains an open challenge.
Acknowledgments
This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Science and Technology Program (Shenzhen Key Laboratory Grant No. ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing.
@@ A.3 Comparison-base
Appendix B Detailed Slot Guessing Pipeline
B.1 Back-Translation
The back-translation function applies a two-step translation process to generate a paraphrased caption $S^{\prime}_{i}$ from the original caption $S_{i}$. In this method, we use the Google Translate API to translate the caption into Chinese and then back into the original language to generate the paraphrase.
Algorithm 1 Back-Translation
1:Input: Original caption $S_{i}$
2: Translate $S_{i}$ to an intermediate language $L$
3: Translate the resulting caption back from language $L$ to the original language
4:Output: Paraphrased caption $S^{\prime}_{i}$
B.2 Keyword Extraction
We extract keywords from both the original caption $S_{i}$ and the paraphrased caption $S^{\prime}_{i}$ using the Stanford POS Tagger. Keywords are identified as nouns (NN), adjectives (JJ), and verbs (VB), which are considered to encapsulate the core meaning of the sentence. We apply this process to both captions.
Algorithm 2 Keyword Extraction
1:Input: Caption $S$
2: Apply POS tagging to $S$ to obtain tags for each word
3: Extract words whose POS tags are in {NN, $J$, VB}
4:Output: List of extracted keywords $K$
B.3 Keyword Masking
We apply a masking function to randomly select one keyword from the extracted keywords and replace it with a placeholder token [MASK]. This is done by identifying the position of the selected keyword in the sentence and substituting it with the placeholder.
Algorithm 3 Keyword Masking
1:Input: Caption $S$, Keywords $K$
2:If $K$ is empty then return "failed"
3: Randomly select a keyword $k$ from $K$
4: Find the first occurrence of $k$ in $S$
5: Replace $k$ with the placeholder [MASK]
6:Output: Masked caption $S_{\text{mask}}$
Appendix C Contamination Degree Analysis
Based on  §4.1, the degrees on multi-choice datasets are defined as: $\Delta\in(-1.6,-0.2]$ for minor leakage, $\Delta\in(-2.9,-1.6]$ for partial leakage, and $\Delta\leq-2.9$ for severe leakage. Based on §4.1, the degrees on caption datasets are defined as: $\Delta\in(-2.4,-1.1]$ for minor leakage, $\Delta\in(-5.0,-2.4]$ for partial leakage, and $\Delta\leq-5.0$ for severe leakage. Details are shown in the algorithm 4.
Algorithm 4 Contamination Degree Analysis
0: Benchmark dataset $D$, Model $M$
2:if $D$ is multiple-choice then
3:  Generate perturbed set $D_{\text{pert}}$ via §3.1
4:else
5:  Generate perturbed set $D_{\text{pert}}$ via §3.2
6:end if
7: Compute CRCR, PCRPCR, $\Delta$, $\Phi$ using §3.3
8:if multiple-choice then
10:else
12:end if
12:CRCR, PCRPCR, $\Delta$, $\Phi$, $\mathcal{C}$
Appendix D Other Experiments
D.1 Semantic & Lexical Similarity After Back-Translation
Setup.
[…]
  • SBERT cosine similarity Reimers and Gurevych (2019) as a sentence-level semantic score, and
  • BLEU-4 Papineni et al. (2002) as a token-overlap lexical score.
We additionally report the Pearson correlation between the two metrics across captions within each dataset.
Dataset Avg. SBERT $\uparrow$ Avg. BLEU $\uparrow$ Correlation $r$
COCO Caption 0.894 0.236 0.386
NoCaps 0.887 0.264 0.410
Vintage 0.914 0.441 0.423
Table 10: Average semantic (SBERT) and lexical (BLEU-4) similarity between original and back-translated captions, together with their Pearson correlation ($r$).
Key Observations.
  • High semantic preservation. All three datasets record SBERT scores close to 0.9, indicating that back-translation keeps the meaning of captions largely intact; the Vintage split achieves the strongest preservation (0.914).
  • Substantial lexical variation. BLEU-4 values are comparatively low, showing that wording and surface forms differ considerably—consistent with the presence of synonym substitutions and syntactic reshuffling introduced by back-translation.
  • Weak yet positive coupling. Pearson correlations between the two metrics lie in the 0.380.38-$0.42$ band, suggesting only a mild positive relationship: captions that keep more tokens also tend to retain semantics, but plenty of cases preserve meaning even with low lexical overlap.
These results justify using back-translation as a semantics-preserving yet lexically diversifying perturbation in our contamination-detection pipeline.
D.2 Sanity Check for the “I don’t know” Instruction
Setup.
To verify that appending the uncertainty clause “If you do not know the answer, output “I don’t know”.” effectively suppresses random guessing, we performed a pilot experiment on 1 000 randomly sampled questions from MMStar. All images were removed, so a truly vision-grounded model should either fail or explicitly abstain. We evaluated the unimodal LLaMA2-7B language model under two settings:
  • Deter: deterministic decoding with the uncertainty instruction;
  • Non-Deter: deterministic decoding without the instruction.
Results.
Table 11 shows that the instruction causes the model to respond “I don’t know” 238 times and reduces apparent accuracy from 44.8% to 25.6% (a drop of 19.2%). This confirms that nearly half of the seemingly correct answers in the uninstructed setting are likely due to lucky guesses rather than genuine reasoning, justifying our decision to include the clause in all main experiments.
Setting Accuracy (%) # “I don’t know”
Deter (+ instruction) 25.6 238
Non‐Deter (- instruction) 44.8 0
Table 11: Effect of the uncertainty instruction on LLaMA2-7B.
“I don’t know” will therefore be treated as an explicit abstention in the main study, ensuring reported accuracies reflect genuine vision‐language capabilities rather than random chance.
D.3 Prompt Consistency
Our main experiments used a fixed prompt template across all evaluations. However, since different prompt wordings may influence model performance, we conducted additional experiments to assess the robustness of our methods. In particular, we tested two representative MLLMs (Claude-3.5-Sonnet and Gemini-1.5-Pro) on ScienceQA and COCOCaption under alternative prompt templates. Results show that while absolute accuracy values fluctuate slightly, the contamination-sensitive metric $\Delta$ remains nearly unchanged, confirming that our method is robust to prompt variations.
MLLMs Prompt Templates CR PCR $\Delta$ $\Phi$
Claude-3.5-Sonnet Prompt 1 (Orig.) 70.3 65.0 -5.3 15.3
Claude-3.5-Sonnet Prompt 2 70.4 (+0.1) 65.0 (+0.0) -5.4 (-0.1) 15.5 (+0.2)
Claude-3.5-Sonnet Prompt 3 70.4 (+0.1) 65.0 (+0.0) -5.4 (-0.1) 15.5 (+0.2)
Gemini-1.5-Pro Prompt 1 (Orig.) 68.5 67.9 -0.6 6.6
Gemini-1.5-Pro Prompt 2 68.6 (+0.1) 68.0 (+0.1) -0.6 (+0.0) 6.5 (-0.1)
Gemini-1.5-Pro Prompt 3 68.6 (+0.1) 68.0 (+0.1) -0.6 (+0.0) 6.5 (-0.1)
Table 12: Prompt consistency test on ScienceQA. $\Delta$ remains stable despite prompt variations.
MLLMs Prompt Templates CR PCR $\Delta$ $\Phi$
Claude-3.5-Sonnet Prompt 1 (Orig.) 53.7 51.0 -2.7 21.8
Claude-3.5-Sonnet Prompt 2 53.9 (+0.2) 51.1 (+0.1) -2.8 (-0.1) 22.0 (+0.2)
Claude-3.5-Sonnet Prompt 3 53.8 (+0.1) 51.1 (+0.1) -2.7 (+0.0) 21.9 (+0.1)
Gemini-1.5-Pro Prompt 1 (Orig.) 57.5 55.3 -2.2 21.6
Gemini-1.5-Pro Prompt 2 57.6 (+0.1) 55.5 (+0.2) -2.1 (+0.1) 21.5 (-0.1)
Gemini-1.5-Pro Prompt 3 57.6 (+0.1) 55.5 (+0.2) -2.1 (+0.1) 21.6 (+0.0)
Table 13: Prompt consistency test on COCOCaption. Results show negligible effect on $\Delta$.
Option Order Sensitivity Test.
The original prompt was formulated as follows:
Prompt 1 (Original)
Please answer the following multichoice question.
Question: {question}
Reply with answer only.
We then compared two variants:
Prompt 2
Please respond to the following multiple-choice question.
Question: {question}
Provide only the answer.
Prompt 3
Answer the multiple-choice question below.
Question: {question}
Reply with your answer only.
Results on ScienceQA are shown in Table 12. Across all prompt templates, $\Delta$ changes by at most $0.1$, highlighting stability.
Slot Guessing for Perturbed Caption.
The original caption-prompt was:
Prompt 1 (Original)
Fill in the [MASK] of the following sentence in one word:
{caption}
Only reply with the word you fill in the [MASK].
We added two reworded prompts:
Prompt 2
Complete the [MASK] in the sentence below with a single word:
{caption}
Respond only with the word you used to replace [MASK].
Prompt 3
Provide one word to fill in the [MASK] in the following sentence:
{caption}
Your reply should only include the word you have selected for [MASK].
Results on COCOCaption (Table 13) again confirm robustness: CR and PCR shift slightly, but $\Delta$ stays virtually unaffected.
Overall, these experiments confirm that our detection framework is robust to natural prompt variations, further supporting its generality.
Appendix E Detailed Dataset Overlap Analysis
It is impractical to quantify overlapping samples: 1) Many models do not release their complete training datasets publicly; instead, they only mention the data sources in their technical reports. 2) Even if we had access to complete training datasets, identifying specific overlapping samples using matching algorithms (such as exact match) remains challenging. This is because the original benchmarks might have undergone data augmentation before being used for model training, and multimodal benchmarks include images, both of which complicate the practical utility of matching algorithms. The feasible approach is manually reviewing the technical reports of these models to verify whether their training data overlaps with benchmarks, as shown in the table 14.
MLLMs Multimodal Alignment/Pretraining Data Supervised Fine-Tuning Data
Phi-3-Vision Alignment Data includes FLD-5B. Not yet released
Open Images is one source of FLD-5B.
Open Images is also a source of Nocaps.
Therefore, there is potential overlap in Nocaps.
VILA No overlap Includes RefCOCO, VQAv2, GQA
MS COCO is a source of RefCOCO, VQAv2.
GQA’s source is Visual Genome Scene Graph, which also includes MS COCO.
COCO Caption’s source is MS COCO, and NoCaps’ source includes COCO Caption.
Therefore, there is potential overlap in COCO Caption and NoCaps.
Idefics2 Alignment Data includes SBU Captions SFT Data includes SBU Captions: potential overlap in COCO Caption and NoCaps.
SBU Captions’ source includes Flickr
COCO Caption’s source includes MS COCO, and MS COCO’s source includes Flickr
NoCaps’ source includes COCO Caption
Therefore, there is potential overlap in COCO Caption and NoCaps.
LLaVA-1.5 Alignment Data includes SBU Captions: COCO Caption and NoCaps with potential overlap. SFT Data includes RefCOCO, VQAv2, GQA: COCO Caption and NoCaps with potential overlap.
Yi-VL Alignment Data includes Flickr, VQAv2, RefCOCO: SFT Data includes GQA: COCO Caption and NoCaps with potential overlap.
COCO Caption and NoCaps with potential overlap.
DeepSeek-VL2 No overlap SFT Data includes Flickr, GQA: COCO Caption and NoCaps with potential overlap.
Qwen-VL-Chat Directly uses COCO Caption in the pretraining stage, Not yet released
therefore there is partial or entire overlap in COCO Caption and NoCaps.
InternVL2 Alignment Data includes COCO Caption: partial or entire overlap in COCO Caption and NoCaps. SFT Data includes ScienceQA, therefore there is partial or entire overlap in ScienceQA.
Table 14: Comparison of MLLMs and Their Data Sources